OpenStudy (anonymous):

(1/4 + 1/X)/(4+x) WITH LIMITAPPROACHING X TO -4

OpenStudy (anonymous):

i cant use a calc

OpenStudy (inkyvoyd):

rewrite (1/4+1/x) so you can divide both sides by (4+x) and get rid of the indeterminant denom

OpenStudy (anonymous):

4/4x + x/4x?

OpenStudy (inkyvoyd):

sure, whatever

OpenStudy (inkyvoyd):

just tell me what you get and I'll tell you if it's right

OpenStudy (anonymous):

i already no the anwre but i dont know how to do it

OpenStudy (jdoe0001):

\(\bf \Large lim_{x\rightarrow -4} \cfrac{ \frac{ \frac{1}{4}+1 }{x} }{ 4+x } \quad ? \)

OpenStudy (anonymous):

dont btoher helpingme i ur trolling me

OpenStudy (anonymous):

no its (1/x)

OpenStudy (inkyvoyd):

I'm not trolling you

OpenStudy (anonymous):

then just get out

OpenStudy (inkyvoyd):

(1/4 + 1/X)/(4+x) = (16(4+4x)/(4+4x)) does that help?

OpenStudy (anonymous):

the anwer isnt 16

OpenStudy (jdoe0001):

ohh shoot... I see so.... \(\bf\Large lim_{x\rightarrow -4} \cfrac{ \frac{1}{4}+\frac{1}{x} } { 4+x } \) right

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

ur toxic inky

OpenStudy (anonymous):

@jdoe0001 can u help me plz?

OpenStudy (inkyvoyd):

Dude, I got a 5 on my AP BC calc test. also, disagreeing with a CAS is not very smart

OpenStudy (anonymous):

the teacher put in -1/16 im in calc ab

OpenStudy (anonymous):

i got -16

OpenStudy (inkyvoyd):

gimme a sec, I needa write down the problem if I want to make anything useful out of it.

OpenStudy (jdoe0001):

\(\bf lim_{x\rightarrow -4} \cfrac{ \frac{1}{4}+\frac{1}{x} } { 4+x }\\ -------------\\ \cfrac{ \frac{1}{4}+\frac{1}{x} } { 4+x } \implies \cfrac{ \frac{x+4}{4x} }{4+x} \implies \cfrac{x+4}{4x} \times \cfrac{1}{4+x}\\ \implies \cfrac{\cancel{4+x}}{4x} \times \cfrac{1}{\cancel{4+x}}\)

OpenStudy (anonymous):

yes but i need to know how they got the answer withough a calc?

OpenStudy (inkyvoyd):

yeah yeah I'll give the medal to that dude

OpenStudy (jdoe0001):

lol

OpenStudy (inkyvoyd):

@morgan34 sign up and click show steps sigh

OpenStudy (inkyvoyd):

I'm too lazy to work out the problem since I don't have any paper with me, so you may as well get help from the other dude.

OpenStudy (anonymous):

how come u didnt put 4 + x/1

OpenStudy (inkyvoyd):

cause you can't algebra

OpenStudy (jdoe0001):

\(\bf\Large { \cfrac{ \frac{x+4}{4x} }{4+x} \implies \cfrac{ \frac{x+4}{4x} }{\frac{4+x}{1}}}\)

OpenStudy (anonymous):

o yeah orgot ty

OpenStudy (anonymous):

bless

OpenStudy (inkyvoyd):

dude, wolfram alpha works better - just get the 14 day pro trial and type in a problem and click show steps

OpenStudy (inkyvoyd):

cause then you can do ur maths homework in 3 minutes at 4AM

OpenStudy (inkyvoyd):

just don't make a habit out of it or you'll get a 1 on the AP test lol

OpenStudy (anonymous):

mine expired

OpenStudy (inkyvoyd):

make a new account and do the same thing - free users can still get 3 free solutions a day

OpenStudy (anonymous):

ok thanks